OpenStudy (anonymous):

evaluate u/z+xy2, for u=20, x=4, y=7 and z=10

OpenStudy (anonymous):

Substitute your values in. I am not sure if that is supposed to x y^2 or not.. If so then: 20/10 + (4)(7)^2

OpenStudy (anonymous):

these are the answers a.786 b.198 c.294 d.900 im still confused

OpenStudy (anonymous):

20/10 + (4)(7)^2 evaluates to 198. We are just plugging in numbers here. Wherever you see a 'u', you plug in '20'. Wherever you see an 'x' you plug in '4', and wherever you see a 'z' you plug in '10'. The expression becomes: 20/10 + 4 * 7^2 = 2 + 4 * 49 = 198